# Monte Carlo Neural PDE Solver for Learning PDEs via Probabilistic Representation

### v4

@@ 4.4.2 Results
Fig. 7 shows the predicted physical field $u$ of a learned MCNP Solver at $t=1.0$ in E1-E4, respectively. Comparing different $\alpha$ values, we observe that the diffusion effect becomes more pronounced as the fractional coefficient $\alpha$ increases. Table 4 presents each method’s performance and computational cost on the 2D fractional diffusion equation. The results reveal that the MCNP Solver has consistently maintained an error rate under 5% for most tasks and performance indicators. It is interesting to note that the precision of the simulation does not vary monotonically with changes in the fractional coefficient $\alpha$. This phenomenon happens because excessively large or small values of $\alpha$ present challenges to the simulation tasks. For smaller $\alpha$ values, the singularities of the fractional operators would become more remarkable. From a microscopic perspective, particles are more likely to undergo significant jumps following a Lévy process. Conversely, the diffusion rate accelerates with larger $\alpha$ values, leading to a more rapid evolution of the entire physical field. It is worth mentioning that despite the non-rectangular geometry in this section, the MCNP Solver can be extended to a mesh-free regime and preserve the boundary conditions automatically through the random walk of particles.
@@ 5 Additional Experimental Results
Table 5: Compared to the Monte Carlo methods on the 1D convection-diffusion equations with varying $\kappa$. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and the inference time for each method.
E5 ($\beta=0.1,\kappa=0.005$) E6 ($\beta=0.1,\kappa=0.01$)
Method $E_{\ell_{2}}$ (%) $E_{\ell_{\infty}}$ (%) $E_{\ell_{2}}$ (%) $E_{\ell_{\infty}}$ (%) Time (S)
MCM 4.612 $\pm$0.098 6.503 $\pm$0.021 7.248 $\pm$0.301 9.035 $\pm$0.171 0.163
MCM+ 1.427 $\pm$0.031 2.052 $\pm$0.036 2.097 $\pm$0.065 2.730 $\pm$0.068 1.373
MCM++ 0.489 $\pm$0.011 0.693 $\pm$0.009 0.737 $\pm$0.022 0.929 $\pm$0.018 13.710
MCNP-10 0.172 $\pm$0.018 0.222 $\pm$0.018 0.200 $\pm$0.012 0.259 $\pm$0.012 0.004
5.1 Comparison with Monte Carlo Method
In this section, we conduct experiments to compare MCNP Solver with Monte Carlo methods on the convection-diffusion equations (E5 and E6) in Section 4.1. The implementation of the Monte Carlo method utilizes the Feynman-Kac law introduced in Section 3.1. We can directly simulate the corresponding SDE due to the linearity of the convection-diffusion equation. We divide the time interval $[0,2]$ into 200 uniform frames and sample 200/2000/20000 particles per grid point when conducting the Monte Carlo methods, denoted as MCM/MCM+/MCM++, respectively.
Table 5 presents a comparison between MCNP-10 and corresponding Monte Carlo methods. The performances of Monte Carlo methods are limited by the number of particles, especially for large heat transfer rate $\kappa$. Sampling a large number of particles can reduce errors, but it also brings severe computation issues, such as GPU memory and inference time. However, the MCNP Solver can approximate the mathematical expectation in Eq. 7 via the PDF of neighbour grid points rather than sampling a large number of particles, as we introduced in Section 3.3.2.
Table 6: Compared to the pseudo-spectral methods on 2D Navier-Stokes equation with varying $\nu$. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and the inference time for baseline methods and MCNP Solver. We unitize NAN to represent the result whose error is larger than 200%.
E3 ($\nu=10^{-4}$) E4 ($\nu=10^{-5}$)
Method $E_{\ell_{2}}$ (%) $E_{\ell_{\infty}}$ (%) $E_{\ell_{2}}$ (%) $E_{\ell_{\infty}}$ (%) Time (S)
PSM-500 NAN NAN NAN NAN 0.350
PSM-1000 0.163 0.394 NAN NAN 0.629
PSM-2000 0.115 0.231 1.653 6.562 1.220
MCNP-20 2.999 $\pm$0.082 5.336 $\pm$0.131 5.776 $\pm$0.085 10.137 $\pm$0.146 0.008
5.2 Comparison with Traditional Eulerian Solver
In this section, we conduct experiments to compare MCNP Solver with traditional pseudo-spectral methods on the Navier-Stokes equations (E3 and E4) in Section 4.3, which is one kind of state-of-the-art numerical methods on periodical PDEs. The implementation of the pseudo-spectral method is in line with [42], which utilizes the Crank–Nicolson scheme for the temporal revolution. We divide the time interval $[0,10]$ into 500/1000/2000 uniform frames and denote them as PSM-500/1000/2000, respectively.
Table 6 presents a comparison between MCNP-20 and PSM-500/1000/2000. The pseudo-spectral methods must refine the time interval to obtain meaningful simulation results, especially for cases with a low viscosity term $\nu$. PSM-2000 achieves the lowest error rate on all tasks but requires 150 times more inference time than neural PDE solvers. In summary, the precision of the MCNP Solver cannot outperform the numerical solvers when ignoring the inference time, which is also a major drawback for other existing DL-based methods [81, 18]. However, the neural PDE solver has a unique application scenario due to its zero-shot inference nature. As discussed in [81], AI-based methods lack precision compared to classical methods while achieving reasonable accuracy and offering great potential for efficient parameter studies.
@@ 5.2 Comparison with Traditional Eulerian Solver
Table 7: Compared to the FNO on 2D Navier-Stokes equation with varying $\nu$. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and the training costs for baseline methods and MCNP Solver. The training time of MCNP Solver reported in this table is the average over E3 and E4.
@@ 5.3 Comparison with Supervised Solver Learning
Table 7 presents a comparison between MCNP-10/20 and three versions of FNO. Compared to the supervised methods, which use pre-simulated fixed data for training, MCNP Solver can sample new initial fields per epoch, thereby increasing the diversity of training data. As a result, MCNP-10/20 can outperform FNO and FNO+. With the increase of training data, FNO++ achieves better results. This is because the training data is generated from high-precision numerical methods, while the unsupervised methods construct the loss function with low spatiotemporal resolution. However, as a trade-off, FNO++ spends 211-510% more total time than MCNP-20/10.
Table 8: Ablation Studies of each component in MCNP Solver. Relative error (%) and training time for each method on the Navier-Stokes equation tasks with $\nu=10^{-4}$ (E3) and $\nu=10^{-5}$ (E4). The training time of each method reported in this table is the average over E3 and E4.
@@ 5.4 Ablation Studies
Table 8 reports the results and training costs. Compared to MCNP Solver with MCNP-H, the results show that using Heun’s method to simulate the SDEs significantly improves the accuracy of MCNP Solver, while incurring only minimal additional computational cost, especially for MCNP-10. Compared to MCNP Solver with MCNP-I, the interpolation trick plays a crucial role in E4 when $\nu=10^{-5}$. As discussed in Section 3.3.3, extremely low diffusion rates can lead to very short-distance diffusion effects, resulting in $\sum_{i}p_{t}(\boldsymbol{\xi}_{p,t}^{d,i})\delta\ll 1$. Therefore, interpolating the original vorticity field becomes essential to ensure that the PDF of grid points satisfies normalization conditions in Eq. 17. Similarly, the interpolation trick can reduce more relative error for MCNP-20 than MCNP-10 in E3 because fine step size can also introduce localized random walks per step.
@@ 6 Conclusion and Discussion
This paper has several limitations: (1) Some PDEs are not suitable for the Feynman-Kac formula and therefore do not fall within the scope of the MCNP Solver, such as third or higher-order PDEs (involving high-order operators like $u_{xxx}$). (2) The accuracy of the MCNP Solver cannot outperform traditional numerical solvers when disregarding inference time, which is also a major drawback for other existing neural solvers [81, 18], as we discussed in Section 3.3.2.
@@ Acknowledgments
This work was in part supported by National Key Research and Development Program of China [2019YFA0709501] and CAS Project for Young Scientists in Basic Research [No. YSBR-034].
@@
The initial states $u(x,0)$ are generated from the functional space $\mathcal{F}_{N}\triangleq\{\sum_{n=1}^{N}a_{n}\sin(2\pi nx):a_{n}\sim\mathbb{U}(0,1)\},$ where $\mathbb{U}(0,1)$ denotes the uniform distribution over $(0,1)$, and $N$ represents the maximum frequency of the functional space. We generate the ground truth with the pseudo-spectral methods with the Runge-Kutta 2 regime. All PDE instances are generated on the spatial grid $1024$, then down-sampled to $64$. The step size is fixed as $10^{-5}$. We generate 200 test data with seed 1 and 200 validation data with seed 2.
@@
The PINO and MCNP Solver use the 1D FNO as the backbone models. We fix the number of layers and widths as 4 and 32, respectively. We choose the best $mode$ in $\{12,16,20\}$ for FNO, respectively. For PINO and MCNP Solver, we utilize Adam to optimize the neural network for 10000 epochs with an initial learning rate $lr$ and decay the learning rate by a factor of 0.5 every 1000 epochs. The batch size is fixed as 200. The learning rate is chosen from the set $\{0.02,0.01,0.005,0.001\}$. Because the pseudo-spectral methods can naturally stratify the boundary conditions, we fix $\lambda$ as 00 in this section. For PI-DeepONet, we choose the network structure in line with the 1D case in [89] and extend the training iterations to 50000 to ensure the convergence of the model. Moreover, we search the hyperparameters $\lambda_{1}$, $\lambda_{2}$ and the width of the neural networks in $\{1,5,10\}$, $\{1,5,10,25,50\}$ and $\{80,100,120\}$, respectively. Because PI-DeepONet-M utilizes an adaptive re-weighting scheme to balance training loss, we do not need to tune the hyper-parameters in PI-DeepONet. All hyperparameters are chosen via the validation set with seed 0.
@@
The initial states $u(x,0)$ are generated from the functional space $\mathcal{F}_{N}\triangleq\{\sum_{n=1}^{N}a_{n}\sin(2\pi nx):a_{n}\sim\mathbb{U}(0,1)\},$ where $\mathbb{U}(0,1)$ denotes the uniform distribution over $(0,1)$, and $N$ represents the maximum frequency of the functional space. We generate the ground truth with the Python package ‘py-pde’ [100]. All PDE instances are generated on the spatial grid $1024$, then down-sampled to $65$. The step size is fixed as $10^{-6}$. We generate 200 test data with seed 1 and 200 validation data with seed 2.
@@
The PINO and MCNP Solver use the 1D FNO as the backbone models. We fix the number of layers and widths as 4 and 32, respectively. We choose the best $mode$ in $\{12,16,20\}$ for FNO, respectively. For PINO and MCNP Solver, we utilize Adam to optimize the neural network for 20000 epochs with an initial learning rate $lr$ and decay the learning rate by a factor of 0.5 every 2000 epochs. The batch size is fixed as 200. The learning rate is chosen from the set $\{0.02,0.01,0.005,0.001\}$. The hyperparameter $\lambda$ in the PINO loss is chosen from the set $\{1,2,5,10,25,50\}$. For PI-DeepONet, we choose the network structure in line with the 1D case in [89] and extend the training iterations to 200000 to ensure the convergence of the model. Moreover, we search the hyperparameters $\lambda_{1}$, $\lambda_{2}$ and the width of the neural networks in $\{1,5,10\}$, $\{1,5,10,25,50\}$ and $\{80,100,120\}$, respectively. Because PI-DeepONet-M utilizes an adaptive re-weighting scheme to balance training loss, we do not need to tune the hyper-parameters in PI-DeepONet. All hyperparameters are chosen via the validation set with seed 0.
@@
The PINO, SP-FNO and MCNP Solver use the 2D FNO as the backbone models. We fix the number of layers and widths as 4 and 36, respectively. We choose the best $mode$ in $\{12,16,20\}$ for FNO, respectively. For PINO, SP-FNO and MCNP Solver, we utilize Adam to optimize the neural network for 20000 epochs with an initial learning rate of $lr$ and decay the learning rate by a factor of 0.8 every 2000 epochs. The batch size is fixed as 10. The learning rate $lr$ is chosen from the set $\{0.02,0.01,0.005,0.001\}$. Because the pseudo-spectral methods can naturally stratify the boundary conditions, we fix the $\lambda$ as 00 in this section. For FNO, we find that a cosine annealing schedule can obtain the best result when training with the supervised regime. We utilize Adam to optimize the neural network for 400/200/100 epochs with the initial learning rate of $lr$ for FNO/FNO+/FNO++, respectively. The learning rate $lr$ is chosen from the set $\{0.02,0.01,0.005,0.001\}$. All hyperparameters are chosen via the validation set with seed 0.
@@
The test data is generated from the fine-grained Monte Carlo methods with $\Delta t=10^{-3}$, which samples $10^{4}$ particles when simulating the SDEs. The Lévy sampling is calculated via the Python package ‘SciPy’ [85]. Because the simulation of a large number of particles is costly, we generate five test data with seed 1 and five validation data with seed 2.
@@
We utilize the DeepONet as a backbone model and train the MCNP Solver in a mesh-free regime. The batch size is fixed as 20. For each training epoch, we uniformly sample 20000 spatiotemporal points in the target domain and construct the loss function accordingly. We fix the number of layers and widths as 5 and 200, respectively. We use ReLU as an activation function. We utilize Adam to optimize the neural network for 20000 epochs with an initial learning rate of $lr$ and decay the learning rate by 0.8 every 2000 epochs. The learning rate $lr$ is chosen from the set $\{0.02,0.01,0.005,0.001\}$. All hyperparameters are chosen via the validation set with seed 0.
@@
In this section, we discuss the choice of the backbone network of MCNP Solver. Firstly, we test three network structures on the 2D Navier-Stokes equation ($\nu=10^{-4}$ with Li forcing), including FNO [38], UNet [71] and MultiWaveleT- (MWT) based model [20]. We divide the time interval $[0,10]$ into ten uniform lattices for each method. Table 9 presents each backbone method’s performance and computational cost on the 2D Navier-Stokes equations. In this paper, our primary contribution lies in the design of the loss function for unsupervised neural PDE-solving training. Consequently, we choose one of the most widely used backbone models, i.e., FNO, for our main experiments.
Table 9: Choice of the backbone network. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and computational costs for varying backbone models.

### v5

@@ 4.4.2 Results
Fig. 7 shows the predicted physical field $u$ of a learned MCNP Solver at $t=1.0$ in E1-E4, respectively. Comparing different $\alpha$ values, we observe that the diffusion effect becomes more pronounced as the fractional coefficient $\alpha$ increases. Table IV presents each method’s performance and computational cost on the 2D fractional diffusion equation. The results reveal that the MCNP Solver has consistently maintained an error rate under 2% for all tasks and performance indicators. It is interesting to note that the precision of the simulation does not vary monotonically with changes in the fractional coefficient $\alpha$. This phenomenon happens because excessively large or small values of $\alpha$ present challenges to the simulation tasks. For smaller $\alpha$ values, the singularities of the fractional operators would become more remarkable. From a microscopic perspective, particles are more likely to undergo significant jumps following a Lévy process. Conversely, the diffusion rate accelerates with larger $\alpha$ values, leading to a more rapid evolution of the entire physical field. It is worth mentioning that despite the non-rectangular geometry in this section, the MCNP Solver can be extended to a mesh-free regime and preserve the boundary conditions automatically through the random walk of particles.
@@ 5 Additional Experimental Results
TABLE V: In comparing the relative $\ell_{2}$ errors (%) between MCNP and other numerical solvers across four types of equations. In column 9, we report the inference time (S), denoted as $T_{\operatorname{GPU}}/T_{\operatorname{CPU}}$, for a single test sample on both GPU and CPU. The inference time for the fractional PDE on the GPU is not included, as its implementation is based on a CPU package. In column 10, we report the training time of MCNP for each experiment.
Equation Solver E1 E2 E3 E4 E5 E6 $T_{\operatorname{GPU}}/T_{\operatorname{CPU}}$ (S) Train Time (H)
Convection- Diffusion MCM 3.374 $\pm$0.047 4.648 $\pm$0.129 7.269 $\pm$0.382 7.269 $\pm$0.382 4.612 $\pm$0.098 7.248 $\pm$0.301 0.163 / 0.236
MCM+ 1.932 $\pm$0.008 1.440 $\pm$0.050 2.144 $\pm$0.045 6.952 $\pm$0.010 1.427 $\pm$0.031 2.097 $\pm$0.065 0.188 / 0.518
MCM++ 1.718 $\pm$0.003 0.476 $\pm$0.013 0.718 $\pm$0.036 6.896 $\pm$0.010 0.489 $\pm$0.011 0.737 $\pm$0.022 0.279 / 3.045
PSM 0.115 1.121 2.640 3.278 2.428 3.686 0.005 / 0.003
PSM+ 0.056 0.312 0.644 0.733 0.588 0.843 0.007 / 0.005
PSM++ 0.043 0.101 0.165 0.184 0.160 0.209 0.014 / 0.010
MCNP 0.090 $\pm$0.002 0.128 $\pm$0.014 0.170 $\pm$0.012 0.187 $\pm$0.007 0.172 $\pm$0.018 0.200 $\pm$0.012 0.002 / 0.009 0.032-0.054
Allen- Cahn MCM 7.430 $\pm$0.102 7.458 $\pm$0.096 16.155 $\pm$0.343 18.133 $\pm$0.459 7.368 $\pm$0.203 12.068 $\pm$0.023 0.048 / 0.257
MCM+ 2.902 $\pm$0.059 2.992 $\pm$0.071 8.772 $\pm$0.570 10.079 $\pm$0.521 3.449 $\pm$0.048 10.821 $\pm$0.068 0.063 / 2.692
MCM++ 2.103 $\pm$0.014 2.129 $\pm$0.028 1.265 $\pm$0.002 1.848 $\pm$0.005 2.590 $\pm$0.022 7.110 $\pm$0.040 0.551 / 14.568
FDM NAN NAN NAN NAN 1.812 12.441 0.005 / 0.002
FDM+ 0.983 1.468 1.256 2.166 0.480 1.987 0.021 / 0.009
FDM++ 0.625 0.788 1.424 1.140 0.366 1.886 0.043 / 0.018
MCNP 0.547 $\pm$0.053 0.831 $\pm$0.200 0.394 $\pm$0.037 0.582 $\pm$0.078 0.285 $\pm$0.010 1.487 $\pm$0.090 0.002 / 0.009 0.133-0.651
Navier- Stokes MCM 7.326 $\pm$0.035 7.452 $\pm$0.045 5.864 $\pm$0.032 4.884 $\pm$0.012 6.890 $\pm$0.026 6.994 $\pm$0.011 0.180 / 4.552
MCM+ 5.396 $\pm$0.043 7.055 $\pm$0.010 4.523 $\pm$0.026 4.648 $\pm$0.007 5.316 $\pm$0.035 6.700 $\pm$0.014 0.225 / 7.968
MCM++ 3.744 $\pm$0.006 6.814 $\pm$0.005 3.453 $\pm$0.008 4.491 $\pm$0.003 4.023 $\pm$0.025 6.519 $\pm$0.011 0.469 / 16.647
PSM 4.583 12.336 6.116 11.518 4.500 9.920 0.069 / 0.107
PSM+ 4.227 11.427 3.508 8.443 4.122 9.335 0.139 / 0.204
PSM++ 4.102 11.124 2.946 6.789 4.007 9.147 0.344 / 0.513
MCNP 2.397 $\pm$0.101 7.055 $\pm$0.123 2.999 $\pm$0.082 5.776 $\pm$0.085 4.609 $\pm$0.048 10.161 $\pm$0.150 0.007 / 0.065 0.252-0.531
Fractional MCM 5.668 $\pm$0.629 5.710 $\pm$0.886 5.483 $\pm$0.866 6.511 $\pm$0.198 – / 0.058
MCM+ 5.139 $\pm$0.219 5.330 $\pm$0.236 4.074 $\pm$0.106 2.821 $\pm$0.168 – / 0.518
MCM++ 5.057 $\pm$0.066 5.129 $\pm$0.056 3.799 $\pm$0.088 2.015 $\pm$0.010 – / 6.625
PSM 0.114 0.499 1.908 4.644 – / 0.011
PSM+ 0.061 0.252 0.928 2.076 – / 0.018
PSM++ 0.032 0.114 0.378 0.820 – / 0.039
MCNP 1.139 $\pm$0.145 1.538 $\pm$0.616 1.424 $\pm$0.266 0.977 $\pm$0.045 – / 0.010 0.139
5.1 Comparison with Numerical Methods
This section presents experiments comparing the MCNP with various numerical solvers, including the traditional Monte Carlo method and Eulerian methods, such as PSM and FDM, for all equations in Section 4. For Monte Carlo methods, we directly simulate the corresponding SDEs for convection-diffusion and fractional equations, using the branching diffusion method [24] for Allen-Cahn and the random vortex method [66] for Navier-Stokes. To evaluate the effect of particle count, we vary the number of particles, with more plus signs (MCM/MCM+/MCM++) indicating a higher number of particles. For Eulerian solvers, we employ FDM for Allen-Cahn equations and PSM for the other three types of PDEs. In this case, we vary the step size, with more plus signs (FDM/FDM+/FDM++ or PSM/PSM+/PSM++) representing a smaller step size. To ensure a fair comparison, we test the solvers across multiple spatial resolutions, selecting the lowest resolution that yields comparable results. The implementation details are provided in Appendix I.5. Table V summarizes the comparison between MCNP and other numerical solvers.
According to the results in Table V, the performance and efficiency of Monte Carlo methods are usually heavily constrained by the number of particles, particularly at high diffusion rates. For instance, when $\kappa$ is large (as in E3 and E6), there is a significant error difference between MCM and MCM++ in the convection-diffusion equation. While increasing the number of particles can reduce errors, it introduces serious computational challenges, including memory and inference time. In contrast, the MCNP solver approximates the expectation in Eq. 7 using the PDF of neighboring grid points, avoiding the need to sample a large number of particles, as discussed in Section 3.3.2.
In comparing MCNP with traditional Eulerian methods, we observe that conventional solvers often require refining the time step, especially in highly nonlinear cases, leading to increased computational costs. In contrast, MCNP offers significant speed advantages, particularly in solving equations like Navier-Stokes equations, achieving 10-50x speedup on GPU and 2-8x on CPU. Nevertheless, the speedup is limited in some 1D experiments, such as E1 for convection-diffusion, with MCNP sometimes underperforming on the CPU. This suggests that traditional methods can still be more efficient in some simple cases, where the computational complexity is reduced. Another interesting observation is that the GPU provides a more significant speedup for MCNP than for traditional solvers. Such greater GPU speedup is due to MCNP’s ability to leverage spatiotemporal parallelism and output physical fields directly at any time step without sequential iteration. On the other hand, traditional methods often struggle to take full advantage of GPU acceleration, as they are typically designed with sequential computations in the temporal dimension.
Furthermore, like other neural operators, the MCNP Solver requires hyperparameter tuning and training, while traditional numerical methods do not. However, once training is complete, the MCNP Solver can perform inference on new initial value problems without additional training or tuning. This feature makes neural operators particularly promising for applications in inverse design and real-time physical simulations, such as weather forecasting and fluid control [83, 37]. In this section, we report the inference time for the MCNP Solver, while the hyperparameter tuning costs are discussed in Appendix IV.
@@ 5.1 Comparison with Numerical Methods
TABLE VI: Compared to the FNO on 2D Navier-Stokes equation with varying $\nu$. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and the training costs for baseline methods and MCNP Solver. The training time of MCNP Solver reported in this table is the average over E3 and E4.
@@ 5.2 Comparison with Supervised Solver Learning
Table VI presents a comparison between MCNP-10/20 and three versions of FNO. Compared to the supervised methods, which use pre-simulated fixed data for training, MCNP Solver can sample new initial fields per epoch, thereby increasing the diversity of training data. As a result, MCNP-10/20 can outperform FNO and FNO+. With the increase of training data, FNO++ achieves better results. This is because the training data is generated from high-precision numerical methods, while the unsupervised methods construct the loss function with low spatiotemporal resolution. However, as a trade-off, FNO++ spends 211-510% more total time than MCNP-20/10.
TABLE VII: Ablation Studies of each component in MCNP Solver. Relative error (%) and training time for each method on the Navier-Stokes equation tasks with $\nu=10^{-4}$ (E3) and $\nu=10^{-5}$ (E4). The training time of each method reported in this table is the average over E3 and E4.
@@ 5.3 Ablation Studies
Table VII reports the results and training costs. Compared to MCNP Solver with MCNP-H, the results show that using Heun’s method to simulate the SDEs significantly improves the accuracy of MCNP Solver, while incurring only minimal additional computational cost, especially for MCNP-10. Compared to MCNP Solver with MCNP-I, the interpolation trick plays a crucial role in E4 when $\nu=10^{-5}$. As discussed in Section 3.3.3, extremely low diffusion rates can lead to very short-distance diffusion effects, resulting in $\sum_{i}p_{t}(\boldsymbol{\xi}_{p,t}^{d,i})\delta\ll 1$. Therefore, interpolating the original vorticity field becomes essential to ensure that the PDF of grid points satisfies normalization conditions in Eq. 17. Similarly, the interpolation trick can reduce more relative error for MCNP-20 than MCNP-10 in E3 because fine step size can also introduce localized random walks per step.
@@ 6 Conclusion and Discussion
This paper has several limitations: (1) Some PDEs are not suitable for the Feynman-Kac formula and therefore do not fall within the scope of the MCNP Solver, such as third or higher-order PDEs (involving high-order operators like $u_{xxx}$). (2) The accuracy of the MCNP Solver cannot outperform traditional numerical solvers when disregarding inference time, which is also a major drawback for other existing neural solvers [83, 19], as we discussed in Section 3.3.2. (3) Like other neural operators, the MCNP Solver requires hyperparameter tuning and training, while traditional numerical methods do not. However, once trained, neural operators can be applied directly to new initial value problems for rapid inference.
@@ Acknowledgments
This work was in part supported by NSFC [No. 9247010235], National Key Research and Development Program of China [2019YFA0709501] and CAS Project for Young Scientists in Basic Research [No. YSBR-034].
@@
The initial states $u(x,0)$ are generated from the functional space $\mathcal{F}_{N}\triangleq\{\sum_{n=1}^{N}a_{n}\sin(2\pi nx):a_{n}\sim\mathbb{U}(0,1)\},$ where $\mathbb{U}(0,1)$ denotes the uniform distribution over $(0,1)$, and $N$ represents the maximum frequency of the functional space. We generate the ground truth with the pseudo-spectral methods with the Crank–Nicolson regime. All PDE instances are generated on the spatial grid $1024$, then down-sampled to $64$. The step size is fixed as $10^{-5}$. We generate 200 test data with seed 1 and 200 validation data with seed 2.
@@
The PINO and MCNP Solver use the 1D FNO as the backbone models. We fix the number of layers and widths as 4 and 32, respectively. We choose the best $modes$ (the number of frequency components) in $\{12,16,20\}$ for FNO, respectively. For PINO and MCNP Solver, we utilize Adam to optimize the neural network for 10000 epochs with an initial learning rate $lr$ and decay the learning rate by a factor of 0.5 every 1000 epochs. The batch size is fixed as 200. The learning rate is chosen from the set $\{0.02,0.01,0.005,0.001\}$. Because the pseudo-spectral methods can naturally stratify the boundary conditions, we fix $\lambda$ as 00 in this section. For PI-DeepONet, we choose the network structure in line with the 1D case in [90] and extend the training iterations to 50000 to ensure the convergence of the model. Moreover, we search the hyperparameters $\lambda_{1}$, $\lambda_{2}$ and the width of the neural networks in $\{1,5,10\}$, $\{1,5,10,25,50\}$ and $\{80,100,120\}$, respectively. Because PI-DeepONet-M utilizes an adaptive re-weighting scheme to balance training loss, we do not need to tune the hyper-parameters in PI-DeepONet. All hyperparameters are chosen via the validation set with seed 0.
@@
The initial states $u(x,0)$ are generated from the functional space $\mathcal{F}_{N}\triangleq\{\sum_{n=1}^{N}a_{n}\sin(2\pi nx):a_{n}\sim\mathbb{U}(0,1)\},$ where $\mathbb{U}(0,1)$ denotes the uniform distribution over $(0,1)$, and $N$ represents the maximum frequency of the functional space. We generate the ground truth with the Python package ‘py-pde’ [102]. The ground truth solution is generated via the finite-difference method with the Runge-Kutta 4 method. All PDE instances are generated on the spatial grid $1024$, then down-sampled to $65$. The step size is fixed as $10^{-6}$. We generate 200 test data with seed 1 and 200 validation data with seed 2.
@@
The PINO and MCNP Solver use the 1D FNO as the backbone models. We fix the number of layers and widths as 4 and 32, respectively. We choose the best $modes$ in $\{12,16,20\}$ for FNO, respectively. For PINO and MCNP Solver, we utilize Adam to optimize the neural network for 20000 epochs with an initial learning rate $lr$ and decay the learning rate by a factor of 0.5 every 2000 epochs. The batch size is fixed as 200. The learning rate is chosen from the set $\{0.02,0.01,0.005,0.001\}$. The hyperparameter $\lambda$ in the PINO loss is chosen from the set $\{1,2,5,10,25,50\}$. For PI-DeepONet, we choose the network structure in line with the 1D case in [90] and extend the training iterations to 200000 to ensure the convergence of the model. Moreover, we search the hyperparameters $\lambda_{1}$, $\lambda_{2}$ and the width of the neural networks in $\{1,5,10\}$, $\{1,5,10,25,50\}$ and $\{80,100,120\}$, respectively. Because PI-DeepONet-M utilizes an adaptive re-weighting scheme to balance training loss, we do not need to tune the hyper-parameters in PI-DeepONet. All hyperparameters are chosen via the validation set with seed 0.
@@
The PINO, SP-FNO and MCNP Solver use the 2D FNO as the backbone models. We fix the number of layers and widths as 4 and 36, respectively. We choose the best $modes$ in $\{12,16,20\}$ for FNO, respectively. For PINO, SP-FNO and MCNP Solver, we utilize Adam to optimize the neural network for 20000 epochs with an initial learning rate of $lr$ and decay the learning rate by a factor of 0.8 every 2000 epochs. The batch size is fixed as 10. The learning rate $lr$ is chosen from the set $\{0.02,0.01,0.005,0.001\}$. Because the pseudo-spectral methods can naturally stratify the boundary conditions, we fix the $\lambda$ as 00 in this section. For FNO, we find that a cosine annealing schedule can obtain the best result when training with the supervised regime. We utilize Adam to optimize the neural network for 400/200/100 epochs with the initial learning rate of $lr$ for FNO/FNO+/FNO++, respectively. The learning rate $lr$ is chosen from the set $\{0.02,0.01,0.005,0.001\}$. All hyperparameters are chosen via the validation set with seed 0.
@@
The test data is generated from a fine-grained pseudo-spectral method, using the Bessel function as the basis function. The radius $[0,1]$ is divided into 400 grid points, with a time step of $10^{-4}$. Afterward, we downsample the spatiotemporal resolution to $40\times 10$. Fifty test datasets are generated using seed 1, and fifty validation datasets are generated using seed 2.
@@
We utilize the DeepONet as a backbone model and train the MCNP Solver in a mesh-free regime. The batch size is fixed as 20. For each training epoch, we uniformly sample 2000 spatiotemporal points in the target domain and construct the loss function accordingly. We fix the number of layers and widths as 5 and 200, respectively. We use ReLU as an activation function. We utilize Adam to optimize the neural network for 10000 epochs with an initial learning rate of $lr$ and decay the learning rate by 0.8 every 2000 epochs. The learning rate $lr$ is chosen from the set $\{0.02,0.01,0.005,0.001\}$. All hyperparameters are chosen via the validation set with seed 0.
I.5: Implementation Details of Numerical Solvers
In this section, we introduce the implementation details of traditional numerical solvers, including both Eulerian and Monte Carlo methods. To ensure a fair comparison, we tested the solvers across multiple spatial resolutions, selecting the lowest resolution that produced comparable results. The implementation of traditional Eulerian methods follows the same procedure as the data generation process; however, to accelerate inference, we employed a coarser spatiotemporal grid. The specific spatiotemporal resolutions used are provided in Table VIII.
TABLE VIII: Spatiotemporal resolutions and number of particles used for testing numerical solvers.
Equation Solvers Temporal Resolution Spatial Resolution Number of Particles
Convection- Diffusion MCM / MCM+ / MCM++ 10 64 200 / 2000 / 20000
PSM / PSM+ / PSM++ 10 / 20 / 50 16
Allen- Cahn MCM / MCM+ / MCM++ 20 64 200 / 2000 / 20000
FDM / FDM+ / FDM++ 20 / 100 / 200 64
Navier- Stokes MCM / MCM+ / MCM++ 10 $64\times 64$ 10 / 20 / 50
PSM / PSM+ / PSM++ 100 / 200 / 500 $16\times 16$
Fractional MCM / MCM+ / MCM++ 20 40 200 / 2000 / 20000
PSM / PSM+ / PSM++ 10 / 20 / 50 20
@@
In this section, we discuss the choice of the backbone network of MCNP Solver. Firstly, we test three network structures on the 2D Navier-Stokes equation ($\nu=10^{-4}$ with Li forcing), including FNO [40], UNet [73] and MultiWaveleT- (MWT) based model [21]. We divide the time interval $[0,10]$ into ten uniform lattices for each method. Table IX presents each backbone method’s performance and computational cost on the 2D Navier-Stokes equations. In this paper, our primary contribution lies in the design of the loss function for unsupervised neural PDE-solving training. Consequently, we choose one of the most widely used backbone models, i.e., FNO, for our main experiments.
TABLE IX: Choice of the backbone network. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and computational costs for varying backbone models.